OpenStudy (anonymous):

Solve for x: (x+4)/3= 2

pooja195 (pooja195):

We flip the fraction an multiply by the reciporcal \[\huge~\rm~\frac{ x+4 }{ 3 }\times (3) = (2) \times(3)\]

OpenStudy (anonymous):

6

pooja195 (pooja195):

good and the 3's cancel out so we get \[\huge~\rm~x+4=6\] from here solve for x

OpenStudy (anonymous):

x=2?

pooja195 (pooja195):

yep :)

OpenStudy (anonymous):

yay :) thnx :)

pooja195 (pooja195):

:)

OpenStudy (anonymous):

x=2

rvc (rvc):

you can recheck the answer by putting the value of x in the equation. if LHS=RHS then its correct :)